please show me all the steps in this problem along with the answer (2/x)^-2
Answer by sdmmadam@yahoo.com(530)

please show me all the steps in this problem along with the answer (2/x)^-2
(2/x)^-2
= 1/[(2/x)^2] using (a)^-m = 1/[(a)^+m]
=1/(2^2/x^2) using (a/b)^m = (a^m)/(b^m)
=1/(4/x^2)
= (x^2)/(4) using 1/[(a/b)] = b/a